OpenStudy (anonymous):

Which of the following is the solution to the equation sqrt5^8n = 125^(n +3)

OpenStudy (anonymous):

Did u mean this? \[\sqrt{5^{8n}} = 125^{n+3}\]

OpenStudy (anonymous):

the 8n is outside of the sqrt function, but the 125^(n+3) you entered looks right

OpenStudy (anonymous):

\[\sqrt{5}^{8n}=125^{n+3}\]like this?

OpenStudy (anonymous):

exactly!

OpenStudy (anonymous):

\[\sqrt{5}^{8n}=125^{n+3}\]

OpenStudy (anonymous):

\[\sqrt{5}^2=5\], \[5^3=125\] so \[\sqrt{5}^6=125\] is a start

OpenStudy (anonymous):

my choices are n = –9 n = 9 n = –3 n = 3

OpenStudy (anonymous):

n = 9

OpenStudy (anonymous):

opps...and also -9

OpenStudy (anonymous):

put \[8n=6(n+3)\] amd solve for n

OpenStudy (anonymous):

n = 9

OpenStudy (anonymous):

Thanks so much!